OpenStudy (fairy733189):

The table below shows four systems of equations: System 1: 4x - 5y =2 3x -y = 8 System 2: 4x - 5y = 2 3x - 2y = 1 System 3: 4x - 5y = 2 3x - 8y = 4 System 4: 4x - 5y = 2 10x - 9y = 4 Which pair of systems will have the same solution? (4 points) System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1 System 2 and system 4, because the second equation in system 4 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 2 and system 4, because the second equation in system 4 is obtained by adding the first equation in system 2 to three times the second equation in system 2

OpenStudy (fairy733189):

@noseboy908

OpenStudy (noseboy908):

I'm here, I'm just working it out for myself first.

OpenStudy (fairy733189):

thats fine!

OpenStudy (noseboy908):

Let's get right in it, then. Recall yesterday how we were discussing every now and then how one moves parts of an equation around?

OpenStudy (fairy733189):

yes

OpenStudy (noseboy908):

Great! So for each set of equations we'll have to choose one of the equations to set it in the form of "x=..", or "y=...", so start by choosing an an equaion from the first set to change to one of those two forms. (Try to pick the one that will be the simplest to move around)

OpenStudy (fairy733189):

I'll choose the second equation

OpenStudy (noseboy908):

Great choice. Now, writing out your steps, tell me what form the second equation can be put in so that x or y is on its own side.

OpenStudy (fairy733189):

3x - y = 8 +y +y 3x = y +8

OpenStudy (noseboy908):

While you can use that method, I'd highly suggest setting y on its own side instead of x. This is because x has a 3 attached to it (via multiplication), and the y is on its own already in a way.

OpenStudy (fairy733189):

so then: 3x - y = 8 -3x -3x -y = 8 - 3x

OpenStudy (noseboy908):

And now finish the last step in simplifying here. There is a '-1' that y is being multiplied by at present.

OpenStudy (noseboy908):

*by y at present

OpenStudy (fairy733189):

so the equation will be y = 8 - 3x?

OpenStudy (noseboy908):

Well, while you removed the -1 from the y, you have to divide the elements of the right side by it as well. (Just flip the signs from subtraction to addition or from addition to subtraction).

OpenStudy (fairy733189):

so then y = 8 - 3x?

OpenStudy (noseboy908):

That's the same equation you just gave me that I said you needed to change.

OpenStudy (fairy733189):

soon I meant y = 8 + 3x

OpenStudy (fairy733189):

*sorry

OpenStudy (noseboy908):

Just one more step! You have to change the 8 too!

OpenStudy (fairy733189):

y = -8 + 3x

OpenStudy (noseboy908):

There ya go. Now we have our revised formula to work with. Now let's take the first formula, which is 4x-5y=2, and plug in what you have equal to y in the above equation into the y variable of the first equation. Understand?

OpenStudy (fairy733189):

I'm not sure but would it be 4x - 5( -8 + 3x) = 2?

OpenStudy (noseboy908):

Exactly! It's great that you know that for this problem. Now, use your skills of equation simplification and simplify this equation as much as possible, showing your work please. :)

OpenStudy (fairy733189):

ok

OpenStudy (fairy733189):

4x - 5(-8 + 3x) = 2 4x - 40 - 15x = 2 +40 +40 4x - 15x = 42

OpenStudy (fairy733189):

I think I did something wrong because I don't know how to do the rest

OpenStudy (fairy733189):

wait I see what I did wrong

OpenStudy (fairy733189):

4x - 5(-8 + 3x) = 2 4x + 40 - 15x = 2 -40 -40 4x - 15x = -38

OpenStudy (fairy733189):

I still don't understand it

OpenStudy (noseboy908):

You are doing fine! Just keep adding and subtracting for now

OpenStudy (fairy733189):

4x - 15x = -38 -11x = -38?

OpenStudy (noseboy908):

As odd as it sounds, that's correct. Now, wait just a moment, as I wish to verify my math for the equations. You could get a drink of preference if you so desire.

OpenStudy (fairy733189):

ok

OpenStudy (noseboy908):

Thanks for the patience. We're at the point where we established that the the first set of equations can be simplified to -11x=38. Now, for the next step, there's one more level of simplification you can go to. See if you can simplify the one more time.

OpenStudy (fairy733189):

x = -3.45?

OpenStudy (noseboy908):

While your math is correct, you are expected to leave it as an improper fraction in simplest form. What is that?

OpenStudy (fairy733189):

38/11?

OpenStudy (fairy733189):

sorry -38/11

OpenStudy (noseboy908):

Your first answer there is correct. So the outcome of the first equation is "x = 38/11". Got it?

OpenStudy (fairy733189):

why wouldn't it be -?

OpenStudy (noseboy908):

You divided a negative number by a negative number. In every circumstance that comes out positive.

OpenStudy (fairy733189):

sorry I didn't realize 38 was a negative

OpenStudy (fairy733189):

but I got it now

OpenStudy (noseboy908):

Great, so let's move on to the second set of equations. What are they?

OpenStudy (fairy733189):

4x - 5y = 2 3x - 2y = 1

OpenStudy (fairy733189):

should I do all the work I did for the first set and show it to you?

OpenStudy (noseboy908):

For the first set?

OpenStudy (fairy733189):

yes

OpenStudy (noseboy908):

Since we've already worked through the first set I feel no need to see it again.

OpenStudy (fairy733189):

ok so let me figure the second set out and i'll give you the answers so you can see if their right.

OpenStudy (noseboy908):

Great! Though please post your work too. And also note that since we solved for x in the first set, you have to ultimately solve for x here as well. (That means solving for y in the first equation you choose)

OpenStudy (fairy733189):

3x -2y = 1 -3x -3x -2y = 1 - 3x which would become 2y = -1 + 3x Is this good so far?

OpenStudy (noseboy908):

Yep, keep going.

OpenStudy (fairy733189):

4x - 5y = 2 4x - 5( -1 + 3x) = 2 4x + 5 - 15x = 2 -5 -5 4x - 15x = -3 -11x = -3 11/3

OpenStudy (noseboy908):

You forgot to divide the elements of the right side by 2 before substituting it into the equation.

OpenStudy (fairy733189):

sorry but what do you mean?

OpenStudy (noseboy908):

At the point where you said "3x -2y = 1 -3x -3x -2y = 1 - 3x which would become 2y = -1 + 3x", you still have to divide EVERYTHING by 2 in order to get y on its own.

OpenStudy (fairy733189):

then it would be y = - 0.5 + 1.5

OpenStudy (noseboy908):

Remember what I said about using irregular fractions instead of decimals.

OpenStudy (fairy733189):

y = - 1/2 + 3/2?

OpenStudy (noseboy908):

There ya go. Now redo your math with that in mind.

OpenStudy (fairy733189):

4x - 5y = 2 4x - 5( -1/2 + 3/2x) = 2 4x + 5/2 - 15/2x = 2 -5/2 -5/2 4x - 15/2x = -1/2 -3/2 = -1/2 Would we simplify this more?

OpenStudy (noseboy908):

Just a minor error was made. go back to the step of "4x - 15/2x = -1/2", and recheck the math you did for the left side.

OpenStudy (fairy733189):

-7/2 = -1/2

OpenStudy (noseboy908):

Yes, just don't forget the fact that it's -7/2(x)! Now, final step here! Solve for x!

OpenStudy (fairy733189):

-7/2 = -1/2 x = 7

OpenStudy (noseboy908):

Try again. Remember to flip the fraction on the left when you multiply it by the fraction on the right.

OpenStudy (fairy733189):

I got the same answer

OpenStudy (fairy733189):

it's -7/2 * 2/1 which equals 7 right?

OpenStudy (noseboy908):

Your "-1/2" magically turned into "2/1"!

OpenStudy (fairy733189):

when you divide by fractions aren't you supposed to flip them?

OpenStudy (fairy733189):

the second one?

OpenStudy (noseboy908):

Only the fraction that is being moved from one side to another gets flipped. This is because you have to flip it to cancel it out on the side it's already on, and you have to do with the other side what you do to the other.

OpenStudy (fairy733189):

so 2/7?

OpenStudy (fairy733189):

ugh sorry 4/7

OpenStudy (noseboy908):